# A mathematical theory of topological invariants of quantum spin systems

### v2

@@ 4.2 A category of fuzzy semilinear sets
Let us show that ${\mathcal{C}\mathcal{S}}_{n}$ has all pairwise meets, using the notion of transverse intersection from the previous section. Recall (Definition 3.5) that we say two sets $U,V\subset{\mathbb{R}}^{n}$ are transverse if for some $C>0$ we have $d(x,U\cap V)\leq C\max(d(x,U),d(x,V))$ for all $x\in{\mathbb{R}}^{n}$.
@@
Lemma 4.4.
Let $P$ and $Q$ be polyhedra ${\mathbb{R}}^{n}$. If $P\cap Q\neq\varnothing$ then $P$ and $Q$ are transverse.
Proposition 4.2.
@@
Let $U=\cup_{i}P_{i}$ and $V=\cup_{j}P_{j}$ be a decomposition of $U$ and $V$ into finite unions of polyhedra and choose $r>0$ so that $P_{i}^{r}\cup Q_{j}^{r}$ is nonempty for each pair $i,j$. By Corollary 4.1 and Lemma 4.4 there are constants $C_{P_{i}Q_{j}}>0$ such that
$d(x,U^{r}\cap V^{r})\leq C\max(d(x,U^{r}),d(x,V^{r}))$
for each pair $i,j$. Since $U^{r}=\cup_{i}P_{i}^{r}$ and $V^{r}=\cup_{i}Q_{i}^{r}$, for any $x\in{\mathbb{R}}$ there are indices $i^{*}$ and $j^{*}$ such that $d(x,U^{r})=d(x,P_{i^{*}}^{r})$ and $d(x,V^{r})=d(x,Q_{j^{*}}^{r})$. Then we have
$d(x,U^{r}\cap V^{r})$ $=d(x,\cup_{ij}(P_{i}^{r}\cap Q_{j}^{r}))$
$\leq d(x,P_{i^{*}}^{r}\cap Q_{j^{*}}^{r}))$
$\leq C_{P_{i^{*}}Q_{j^{*}}}\max(d(x,P_{i^{*}}),d(x,Q_{j^{*}}))$
$=C_{P_{i^{*}}Q_{j^{*}}}\max(d(x,U),d(x,V)),$
@@
Since $U$ and $U^{r}$ (resp. $V$ and $V^{r}$) are isomorphic in ${\mathcal{C}\mathcal{S}}_{n}$, it suffices to show that $U^{r}\cap V^{r}$ is a meet of $U^{r}$ and $V^{r}$. It’s clear that $U^{r}\cap V^{r}\leq U^{r}$ and $U^{r}\cap V^{r}\leq V^{r}$. Now suppose $W\in{\mathcal{C}\mathcal{S}}_{n}$ satisfies $W\leq U^{r}$ and $W\leq V^{r}$. Then there is an $s>0$ such that every $x\in W$ satisfies $\max(d(x,U^{r}),d(x,V^{r}))\leq s$. Since $d(x,U^{r}\cap V^{r})\leq C\max(d(x,U^{r}),d(x,V^{r}))$ for some $C>0$ we have $W\subset(U^{r}\cap V^{r})^{Cs}$. ∎
Proposition 4.3.
@@
We need to show that for any $U,V,W\in{\mathcal{C}\mathcal{S}}_{n}$ we have $U\wedge(V\vee W)\leq(U\wedge V)\vee(U\wedge W)$. According to Corollary 4.2, there exists $r\geq 0$ such that $U\wedge(V\vee W)\simeq U^{r}\cap(V\cup W)^{r}$. Since $(V\cup W)^{r}=V^{r}\cup W^{r},$ we also have $U\wedge(V\vee W)\simeq(U^{r}\cap V^{r})\cup(U^{r}\cap W^{r})$. On the other hand, $U\wedge V\simeq U^{r}\wedge V^{r}\simeq(U^{r})^{s}\cap(V^{r})^{s}$ for some $s\geq 0$, and $U\wedge W\simeq U^{r}\wedge W^{r}\simeq(U^{r})^{t}\cap(W^{r})^{t}$ for some $t\geq 0$. Thus $(U\wedge V)\vee(U\wedge W)\simeq\left((U^{r})^{s}\cap(V^{r})^{s}\right)\cup\left((U^{r})^{t}\cap(W^{r})^{t}\right)$ for some $s,t\geq 0$. Since we have inclusions $U^{r}\cap V^{r}\subseteq(U^{r})^{s}\cap(V^{r})^{s}$ and $U^{r}\cap W^{r}\subseteq(U^{r})^{t}\cap(W^{r})^{t}$, the lemma is proved. ∎
@@
More generally, for any $W\in{\mathcal{C}\mathcal{S}}_{n}$ we may consider a full sub-category ${{{\mathcal{C}\mathcal{S}}_{n}}/W}$ whose objects are $U\in{\mathcal{C}\mathcal{S}}_{n}$ such that $U\leq W$. This is a distributive pre-ordered set, and we will also have occasion to consider local Lie algebras on the associated site.
@@ 4.3 Spherical CS sets
A cone in ${\mathbb{R}}^{n}$ is a non-empty subset of ${\mathbb{R}}^{n}$ which is invariant under $x\mapsto\lambda x$, where $\lambda\geq 0$. Every cone contains the origin 00. Cones in ${\mathbb{R}}^{n}$ are in bijection with subsets of $S^{n-1}=({\mathbb{R}}^{n}\backslash\{0\})/{\mathbb{R}}^{*}_{+}$ where ${\mathbb{R}}^{*}_{+}$ is the group of positive real numbers under multiplication. If $A\subset S^{n-1}$, we denote the corresponding cone $c(A)$. In particular, $c(\varnothing)=\{0\}\in{\mathbb{R}}^{n}$. If $K$ is a cone in ${\mathbb{R}}^{n}$, we will denote the corresponding subset of $S^{n-1}$ by $\hat{K}$.
Definition 4.1.
$A\subset S^{n-1}$ is a spherical polyhedron iff $c(A)$ is a polyhedron and $A$ is contained in some open hemisphere of $S^{n-1}$. $A\subseteq S^{n-1}$ is a spherical CS set iff it is a union of a finite number of spherical polyhedra. The set of spherical CS sets in $S^{n-1}$ is denoted ${\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$.
Every polyhedron is convex and thus contractible. This implies:
Proposition 4.4.
Any spherical polyhedron is contractible.
@@
Without loss of generality, we may assume that the spherical polyhedron is contained in the hemisphere $S^{n-1}_{+}=S^{n-1}\cap\{x_{n}>0\}$. The map ${\mathbb{R}}^{n-1}\rightarrow S^{n-1}_{+}$ which sends $(x_{1},\ldots,x_{n-1})$ to the equivalence class of $(x_{1},\ldots,x_{n-1},1)$ is a homeomorphism which establishes a bijection between bounded polyhedra in ${\mathbb{R}}^{n-1}$ and spherical polyhedra contained in $S^{n-1}_{+}$. ∎
Proposition 4.5.
@@
Clear from definitions. ∎
@@
We proceed first by defining a functor from ${\mathcal{C}\mathcal{S}}_{n}$ to ${\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$. Notice, it is sufficient to define a functor on isomorphism classes of polyhedrons then extend by joins. Every closed polyhedron is a finite intersection of closed half spaces $\bigcap_{i=1}^{k}\{n_{i}\cdot x\leq b_{i}\},$ where for each $i=1,\ldots k$, $n_{i}\in{\mathbb{R}}^{n}$ is the unit normal vector of the $i$th supporting hyperplane and $b_{i}\in{\mathbb{R}}$. By definition, $\bigcap_{i=1}^{k}\{n_{i}\cdot x\leq b_{i}\}$ is isomorphic to $\bigcap_{i=1}^{k}\{n_{i}\cdot x\leq 0\}$ in ${\mathcal{C}\mathcal{S}}_{n}$. In other words, every polyhedron is isomorphic to a cone $\{Nx\preceq 0\},$ where $N$ is the matrix with $n_{i}$ as rows and we write $y\preceq 0$ for $y\in{\mathbb{R}}^{n}$ when $y_{i}\leq 0$ for all $i=1,\ldots n$. If $\{n_{i}\}_{i=1}^{k}$ spans ${\mathbb{R}}^{n}$, the set $\{Nx\preceq 0\}$, which contains no antipodal points, is mapped to its corresponding spherical polyhedron. Otherwise, complete $\{n_{i}\}_{i=1}^{k}$ into a spanning set $\{n_{1},\dots,n_{k},m_{1},\dots,m_{l}\}$ and decompose $\{Nx\preceq 0\}$ into a union of cones $\{Nx\preceq 0\}\cap\bigcap_{j}\{\pm m_{j}\cdot x\leq 0\}$. By the universal property of the join, an arbitrary closed semilinear set is isomorphic to a union of conical polyhedra. For functoriality, it is sufficient to note that for any pair of closed cones $U,V$ a fuzzy inclusion $U\subseteq V^{r}$ for some $r\geq 0$ implies $U\subseteq V$. It is easy to check that this functor is fully faithful and (essentially) surjective. ∎
@@ 4.4 Spherical CS cohomology
5 Local Lie algebras over fuzzy semilinear sets
@@ 5.1 Basic examples
Let $\psi$ be a state of a quantum lattice system on ${\mathbb{R}}^{n}$. By Lemma 3.2 the maps sending $U\in{\mathcal{C}\mathcal{S}}_{n}$ to ${\mathfrak{D}}_{al}(U)$ and ${\mathfrak{D}}_{al}^{\psi}(U)$ are pre-cosheaves of Fréchet spaces on ${\mathcal{C}\mathcal{S}}_{n}$. We denote these pre-cosheaves by ${\mathfrak{D}}_{al}$ and ${\mathfrak{D}}_{al}^{\psi}$. Putting together Lemma 2.1, Proposition 3.5, Corollary 3.1, and Proposition 3.8, we have
Theorem 5.1.
${\mathfrak{D}}_{al}$ is a local Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$. If $\psi$ is gapped, then ${\mathfrak{D}}_{al}^{\psi}$ is a local Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$.
Our main object of study is the local Lie algebra ${\mathfrak{D}}^{\psi}_{al}$ attached to a gapped state $\psi$ of a lattice system $(\Lambda,\{V_{j}\}_{j\in\Lambda})$.
A much simpler example of a local Lie algebra arises from a finite-dimensional Lie algebra ${\mathfrak{g}}$ and any subset $\Lambda\subset{\mathbb{R}}^{n}$. For any $U\in{\mathcal{C}\mathcal{S}}_{n}$ let ${{\mathfrak{g}}_{al}}(U)$ be the space of bounded functions $U\cap\Lambda\rightarrow{\mathfrak{g}}$ which decay superpolynomially away from $U$. The subscript “al” stands for “almost localized”. More precisely, ${{\mathfrak{g}}_{al}}({\mathbb{R}}^{n})$ is the space of bounded functions $\Lambda\rightarrow{\mathfrak{g}}$, while ${{\mathfrak{g}}_{al}}(U)$ is defined as a subspace of ${{\mathfrak{g}}_{al}}({\mathbb{R}}^{n})$ consisting of functions $f:\Lambda\rightarrow{\mathfrak{g}}$ such that the following seminorms are finite:
$p_{k,U}(f)=\sup_{j\in\Lambda}|f(j)|(1+d(U,j))^{k},\quad k\in{\mathbb{N}}.$
Proposition 5.1.
The assignment $U\mapsto{{\mathfrak{g}}_{al}}(U)$ is a local Lie algebra.
@@
It is easy to check that the assignment $U\mapsto{{\mathfrak{g}}_{al}}(U)$ it is a coflasque pre-cosheaf of Fréchet-Lie algebras satisfying Property I. The only thing left to check is that it is a cosheaf of vector spaces. By Lemma 2.1, it is sufficient to show that for any $U,V\in{\mathcal{C}\mathcal{S}}_{n}$ the sequence
${{\mathfrak{g}}_{al}}(U\wedge V)\rightarrow{{\mathfrak{g}}_{al}}(U)\oplus{{\mathfrak{g}}_{al}}(V)\rightarrow{{\mathfrak{g}}_{al}}(U\vee V)\rightarrow 0$
@@
$f_{U}(x)=\begin{cases}f(x),&d(x,U)<d(x,V),\\ \frac{1}{2}f(x),&d(x,U)=d(x,V),\\ 0,&d(x,U)>d(x,V),\end{cases}$
@@
To show exactness at ${{\mathfrak{g}}_{al}}(U)\oplus{{\mathfrak{g}}_{al}}(V)$, suppose $f\in{{\mathfrak{g}}_{al}}(U)\cap{{\mathfrak{g}}_{al}}(V)$. From the proof of Prop. 4.2, there exist $r\geq 0,C_{UV}>0$ such that $d(x,U^{r}\cap V^{r})\leq C_{UV}\max(d(x,U),d(x,V))$. We may assume that $C_{UV}\geq 1$, in which case for any $x\in{\mathbb{R}}^{n}$ and any $k\in{\mathbb{N}}$
$(1+d(x,U^{r}\cap V^{r}))^{k}\leq C^{k}_{UV}(1+\max(d(x,U),d(x,V))^{k}.$
@@
$p_{k,U^{r}\cap V^{r}}(f)\leq C^{k}_{UV}\max(p_{k,U}(f),p_{k,V}(f)).$
Since by Corollary 4.2 one can take $U\wedge V=U^{r}\cap V^{r}$, we conclude that $f\in{{\mathfrak{g}}_{al}}(U\wedge V)$. ∎

### v3

@@ 4.2 A category of fuzzy semilinear sets
Let us show that ${\mathcal{C}\mathcal{S}}_{n}$ has all pairwise meets, using the notion of transverse intersection from the previous section. Recall (Definition 26) that we say two sets $U,V\subseteq{\mathbb{R}}^{n}$ are transverse if for some $C>0$ we have $d(x,U\cap V)\leq C\max(d(x,U),d(x,V))$ for all $x\in{\mathbb{R}}^{n}$.
@@
Lemma 70.
Let $P$ and $Q$ be polyhedra in ${\mathbb{R}}^{n}$. If $P\cap Q\neq\varnothing$ then $P$ and $Q$ are transverse.
Proposition 71.
@@
Let $U=\cup_{i}P_{i}$ and $V=\cup_{j}Q_{j}$ be a decomposition of $U$ and $V$ into finite unions of polyhedra and choose $r>0$ so that $P_{i}^{r}\cap Q_{j}^{r}$ is nonempty for each pair $i,j$. By Corollary 68 and Lemma 70 there are constants $C_{P_{i}Q_{j}}>0$ such that
$d(x,P_{i}^{r}\cap Q_{j}^{r})\leq C_{P_{i}Q_{j}}\max(d(x,P_{i}^{r}),d(x,Q_{j}^{r}))$ (107)
for each pair $i,j$. Since $U^{r}=\cup_{i}P_{i}^{r}$ and $V^{r}=\cup_{j}Q_{j}^{r}$, for any $x\in{\mathbb{R}}^{n}$ there are indices $i^{*}$ and $j^{*}$ such that $d(x,U^{r})=d(x,P_{i^{*}}^{r})$ and $d(x,V^{r})=d(x,Q_{j^{*}}^{r})$. Then we have
$d(x,U^{r}\cap V^{r})$ $=d(x,\cup_{i,j}(P_{i}^{r}\cap Q_{j}^{r}))$
$\leq d(x,P_{i^{*}}^{r}\cap Q_{j^{*}}^{r})$
$\leq C_{P_{i^{*}}Q_{j^{*}}}\max(d(x,P_{i^{*}}^{r}),d(x,Q_{j^{*}}^{r}))$
$=C_{P_{i^{*}}Q_{j^{*}}}\max(d(x,U^{r}),d(x,V^{r})),$ (108)
@@
Since $U$ and $U^{r}$ (resp. $V$ and $V^{r}$) are isomorphic in ${\mathcal{C}\mathcal{S}}_{n}$, it suffices to show that $U^{r}\cap V^{r}$ is a meet of $U^{r}$ and $V^{r}$. It’s clear that $U^{r}\cap V^{r}\leq U^{r}$ and $U^{r}\cap V^{r}\leq V^{r}$. Now suppose $W\in{\mathcal{C}\mathcal{S}}_{n}$ satisfies $W\leq U^{r}$ and $W\leq V^{r}$. Then there is an $s>0$ such that every $x\in W$ satisfies $\max(d(x,U^{r}),d(x,V^{r}))\leq s$. Since $d(x,U^{r}\cap V^{r})\leq C\max(d(x,U^{r}),d(x,V^{r}))$ for some $C>0$ we have $W\subseteq(U^{r}\cap V^{r})^{Cs}$. ∎
Proposition 73.
@@
We need to show that for any $U,V,W\in{\mathcal{C}\mathcal{S}}_{n}$ we have $U\wedge(V\vee W)\leq(U\wedge V)\vee(U\wedge W)$. According to Corollary 72, there exists $r\geq 0$ such that $U\wedge(V\vee W)\simeq U^{r}\cap(V\cup W)^{r}$. Since $(V\cup W)^{r}=V^{r}\cup W^{r},$ we also have $U\wedge(V\vee W)\simeq(U^{r}\cap V^{r})\cup(U^{r}\cap W^{r})$. On the other hand, $U\wedge V\simeq U^{r}\wedge V^{r}\simeq(U^{r})^{s}\cap(V^{r})^{s}$ for some $s\geq 0$, and $U\wedge W\simeq U^{r}\wedge W^{r}\simeq(U^{r})^{t}\cap(W^{r})^{t}$ for some $t\geq 0$. Thus, $(U\wedge V)\vee(U\wedge W)\simeq\left((U^{r})^{s}\cap(V^{r})^{s}\right)\cup\left((U^{r})^{t}\cap(W^{r})^{t}\right)$ for some $s,t\geq 0$. Since we have inclusions $U^{r}\cap V^{r}\subseteq(U^{r})^{s}\cap(V^{r})^{s}$ and $U^{r}\cap W^{r}\subseteq(U^{r})^{t}\cap(W^{r})^{t}$, the proposition is proved. ∎
@@
Definition 74.
The site ${\mathcal{C}\mathcal{S}}_{n}$ has as its objects nonempty closed semilinear sets of ${\mathbb{R}}^{n}$. For any two objects $U,V$, $Mor(U,V)$ contains a single element if $U\subseteq V^{r}$ for some $r\geq 0$ and is empty otherwise. In the former case we write $U\leq V.$ A cover of an object $V$ is a finite collection of objects $U_{i}$, $i\in I$, with $I$ nonempty, such that $U_{i}\leq V$ for all $i\in I$ and $V\leq\bigvee_{i\in I}U_{i}$.
More generally, for any $W\in{\mathcal{C}\mathcal{S}}_{n}$ we may consider a full sub-category ${{\mathcal{C}\mathcal{S}}_{n}}/W$} whose objects are $U\in{\mathcal{C}\mathcal{S}}_{n}$ such that $U\leq W$. This is a distributive pre-ordered set, and we will also have occasion to consider local Lie systems on the associated site.
@@ 4.3 Spherical CS sets
[…]
Definition 75.
$A\subseteq S^{n-1}$ is a spherical polyhedron iff $c(A)$ is a polyhedron and $A$ is contained in some open hemisphere of $S^{n-1}$. $A\subseteq S^{n-1}$ is a spherical CS set iff it is a union of a finite number of spherical polyhedra. The set of spherical CS sets in $S^{n-1}$ is denoted ${\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$.
Remark 76.
For $K$ a polyhedral cone, $\widehat{K}$ need not be contained in an open hemisphere and thus may not be a spherical polyhedron. However, by intersecting $K$ with orthants, one can present every such $K$ as a finite union of polyhedral cones $K_{i}$ such that $\widehat{K}_{i}$ is contained in an open hemisphere. Therefore $\widehat{K}\in{\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$ for every polyhedral cone $K$.
Every nonempty polyhedron is convex and thus contractible. This implies:
Proposition 77.
Any nonempty spherical polyhedron is contractible.
@@
Without loss of generality, we may assume that the spherical polyhedron is contained in the hemisphere $S^{n-1}_{+}=S^{n-1}\cap\{x_{n}>0\}$. The map ${\mathbb{R}}^{n-1}\rightarrow S^{n-1}_{+}$ which sends $(x_{1},\ldots,x_{n-1})$ to the equivalence class of $(x_{1},\ldots,x_{n-1},1)$ is a homeomorphism which establishes a bijection between nonempty bounded polyhedra in ${\mathbb{R}}^{n-1}$ and nonempty spherical polyhedra contained in $S^{n-1}_{+}$. ∎
Proposition 78.
@@
Clear from the definitions. ∎
@@
To prove this proposition, we need two definitions and three lemmas.
Definition 80.
Let $P$ be a nonempty polyhedron in ${\mathbb{R}}^{n}$. Its recession cone is the set
${\rm rec}(P):=\{v\in{\mathbb{R}}^{n}:P+tv\subseteq P\textit{ for all }t\geq 0\}.$ (109)
Lemma 81.
If $P=\bigcap_{i=1}^{k}\{n_{i}\cdot x\leq b_{i}\},$ then ${\rm rec}(P)=\bigcap_{i=1}^{k}\{n_{i}\cdot x\leq 0\}.$ In particular, for any non-empty polyhedron $P$, ${\rm rec}(P)$ is a polyhedral cone, and $\widehat{{\rm rec}(P)}\in{\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$.
@@
Straightforward check. ∎
If $K$ is a polyhedral cone, then clearly ${\rm rec}(K)=K$. More generally, we have
Lemma 82.
For any non-empty polyhedron $P$, $P\simeq{\rm rec}(P)$ as objects in ${\mathcal{C}\mathcal{S}}_{n}$.
Proof.
By the Minkowski-Weyl theorem (see e.g. [33]), $P=Q+{\rm rec}(P)$ where $Q$ is a bounded polyhedron. Clearly, $P\subseteq{\rm rec}(P)^{r}$ where $r=\sup_{q\in Q}\|q\|_{\infty}$. On the other hand, by the definition of ${\rm rec}(P)$ we have $x_{0}+{\rm rec}(P)\subseteq P$ for any $x_{0}\in P$, hence ${\rm rec}(P)\subseteq P^{s}$ where $s=\|x_{0}\|_{\infty}$. ∎
Definition 83.
For $U\in{\mathcal{C}\mathcal{S}}_{n}$, let $\widehat{U}\subseteq S^{n-1}$ be the set of all $u\in S^{n-1}$ such that $c(\{u\})\leq U$.
If $U$ is a polyhedral cone $K$, this definition agrees with the previous definition of $\widehat{K}$. This will be shown in the course of the proof of the next lemma.
Lemma 84.
For any $U\in{\mathcal{C}\mathcal{S}}_{n}$, $\widehat{U}\in{\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$. More precisely, if $U=\cup_{i=1}^{m}P_{i}$ where each $P_{i}$ is a polyhedron, then
$\widehat{U}=\bigcup_{i=1}^{m}\widehat{{\rm rec}(P_{i})}$ (110)
Proof.
First, let us prove (110). If $u\in\widehat{{\rm rec}(P_{i})}$ for some $i$, then $c(\{u\})\leq{\rm rec}(P_{i})$. Since ${\rm rec}(P_{i})\leq P_{i}$ by Lemma 82, this implies $c(\{u\})\leq P_{i}\leq U$, hence $u\in\widehat{U}$.
In the opposite direction, suppose $u\in\widehat{U}$, i.e. $c(\{u\})\subseteq U^{r}$ for some $r\geq 0$. We claim there exists $i_{*}\in\{1,\ldots,m\}$ such that $c(\{u\})\leq P_{i_{*}}$. Indeed, let
$T_{i}=\left\{t\in[0,\infty)|d(tu,P_{i})\leq r\right\}.$ (111)
Here we identified $u\in S^{n-1}$ with the corresponding unit vector in ${\mathbb{R}}^{n}$. We have $\cup_{i}T_{i}=[0,\infty)$, hence there exists $i_{*}\in\{1,\ldots,m\}$ such that $T_{i_{*}}$ is unbounded. Suppose $P_{i_{*}}=\cap_{\alpha=1}^{M}\{n_{\alpha}\cdot x\leq b_{\alpha}\}$ for some vectors $n_{\alpha}\in{\mathbb{R}}^{n}$ and some $b_{\alpha}\in{\mathbb{R}}$. Choose a sequence $\{t_{k}\}_{k\in{\mathbb{N}}}\subseteq T_{i_{*}}$ going to infinity and for each $k\in{\mathbb{N}}$ a point $x_{k}\in P_{i_{*}}$ such that $d(t_{k}u,x_{k})\leq r$. Then for every $\alpha\in\{1,\ldots,M\}$
$t_{k}n_{\alpha}\cdot u\leq b_{\alpha}+n_{\alpha}\cdot(t_{k}u-x_{k})\leq b_{\alpha}+r\|n_{\alpha}\|_{1}.$ (112)
Taking the limit $k\rightarrow\infty$ we conclude that $n_{\alpha}\cdot u\leq 0$ for all $\alpha$, hence by Lemma 81 $c(\{u\})\subseteq{\rm rec}(P_{i_{*}})$ and by Lemma 82 $c(\{u\})\leq P_{i_{*}}$.
The same argument applied to the case when $U$ is a polyhedral cone $K$ shows that the new definition of $\widehat{K}$ agrees with the earlier one.
Finally, since each ${\rm rec}(P_{i})$ is a polyhedral cone, Remark 76 implies that $\widehat{U}\in{\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$. ∎
Proof of Proposition 79.
The map $U\mapsto\widehat{U}$ is order-preserving. Indeed, if $U\leq V$, then $\forall u\in\widehat{U}$ we have $c(\{u\})\leq U\leq V$, hence $u\in\widehat{V}$. Thus $U\mapsto\widehat{U}$ defines a functor from ${\mathcal{C}\mathcal{S}}_{n}$ to ${\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$ which we call $F$. The map $A\mapsto c(A)$ is obviously order-preserving and defines a functor from ${\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$ to ${\mathcal{C}\mathcal{S}}_{n}$ which we call $G$.
We claim that these two functors establish an equivalence between ${\mathcal{C}\mathcal{S}}_{n}$ and ${\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$. Indeed, $F\circ G={\rm id}_{{\mathcal{S}\mathcal{C}\mathcal{S}}_{n}}$ since $\widehat{c(A)}=A$ for any $A\in{\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$. In the opposite direction, let $U=\cup_{i=1}^{m}P_{i}$ where all $P_{i}$ are polyhedra. Then by Lemma 84,
$G\circ F(U)=c\left(\cup_{i=1}^{m}\widehat{{\rm rec}(P_{i})}\right)=\cup_{i=1}^{m}c\left(\widehat{{\rm rec}(P_{i})}\right)=\cup_{i=1}^{m}{\rm rec}(P_{i}).$ (113)
Since by Lemma 82 ${\rm rec}(P_{i})\simeq P_{i}$, this implies $G\circ F(U)\simeq U$. ∎
@@ 4.4 Spherical CS cohomology
5 Local Lie systems over fuzzy semilinear sets
In this section, we define several examples of local Lie systems [^11] In an earlier version, we used the term “local Lie algebras,” which conflicts with a different notion bearing the same name in [14]. over the site ${\mathcal{C}\mathcal{S}}_{n}$. The motivating examples arise from the Lie algebras ${\mathfrak{D}}_{al}(U)$ and ${\mathfrak{D}}^{\psi}_{al}(U)$ studied in Section 2. When $\psi$ is a gapped state invariant under a compact Lie group $G$, we define a pointed graded local Lie system which can be regarded as an equivariant version of ${\mathfrak{D}}^{\psi}_{al}$. Applying the Čech functor gives a pointed DGLA which encodes the properties of the state. Obstructions for the existence of morphisms of pointed DGLAs give rise to topological invariants of $G$-invariant gapped states.
@@ 5.1 Basic examples
Let $\psi$ be a state of a quantum lattice system on ${\mathbb{R}}^{n}$. By Lemma 24 the maps sending $U\in{\mathcal{C}\mathcal{S}}_{n}$ to ${\mathfrak{D}}_{al}(U)$ and ${\mathfrak{D}}^{\psi}_{al}(U)$ are pre-cosheaves of Fréchet spaces on ${\mathcal{C}\mathcal{S}}_{n}$. We denote these pre-cosheaves by ${\mathfrak{D}}_{al}$ and ${\mathfrak{D}}^{\psi}_{al}$. Putting together Lemma 58, Proposition 29, Corollary 34, and Proposition 50, we have
Theorem 87.
${\mathfrak{D}}_{al}$ is a local Lie system over ${\mathcal{C}\mathcal{S}}_{n}$. If $\psi$ is gapped, then ${\mathfrak{D}}^{\psi}_{al}$ is a local Lie system over ${\mathcal{C}\mathcal{S}}_{n}$.
We may also consider quantum lattice systems on subsets of ${\mathbb{R}}^{n}$.
Definition 88.
Let $W\in{\mathcal{C}\mathcal{S}}_{n}$. A quantum lattice system on $W$ is a quantum lattice system on ${\mathbb{R}}^{n}$ such that the lattice $\Lambda$ is contained in $W^{r}$ for some $r\geq 0$.
Obviously, for a quantum lattice system on $W$, ${\mathfrak{D}}_{al}$ is a local Lie system over ${{\mathcal{C}\mathcal{S}}_{n}}/W$}, and for any gapped state $\psi$ of such a system, ${\mathfrak{D}}^{\psi}_{al}$ is a local Lie system on ${{\mathcal{C}\mathcal{S}}_{n}}/W$}.
A much more elementary example of a local Lie system arises from a finite-dimensional Lie algebra $\mathfrak{g}$} and any subset $\Lambda\subseteq{\mathbb{R}}^{n}$. For any $U\in{\mathcal{C}\mathcal{S}}_{n}$ let ${{\mathfrak{g}}_{al}}(U)$ be the space of bounded functions $\Lambda\rightarrow{\mathfrak{g}}$ which decay superpolynomially away from $U$. The subscript “al” stands for “almost localized”. More precisely, ${{\mathfrak{g}}_{al}}({\mathbb{R}}^{n})$ is the space of bounded functions $\Lambda\rightarrow{\mathfrak{g}}$, while ${{\mathfrak{g}}_{al}}(U)$ is defined as a subspace of ${{\mathfrak{g}}_{al}}({\mathbb{R}}^{n})$ consisting of functions $f:\Lambda\rightarrow{\mathfrak{g}}$ such that the following seminorms are finite:
$p_{k,U}(f)=\sup_{x\in\Lambda}|f(x)|(1+d(U,x))^{k},\quad k\in{\mathbb{N}}.$ (114)
Proposition 89.
The assignment $U\mapsto{{\mathfrak{g}}_{al}}(U)$ is a local Lie system.
@@
It is easy to check that the assignment $U\mapsto{{\mathfrak{g}}_{al}}(U)$ is a coflasque pre-cosheaf of Fréchet-Lie algebras satisfying Property I. The only thing left to check is that it is a cosheaf of vector spaces. By Lemma 58, it is sufficient to show that for any $U,V\in{\mathcal{C}\mathcal{S}}_{n}$ the sequence
${{\mathfrak{g}}_{al}}(U\wedge V)\rightarrow{{\mathfrak{g}}_{al}}(U)\oplus{{\mathfrak{g}}_{al}}(V)\rightarrow{{\mathfrak{g}}_{al}}(U\vee V)\rightarrow 0$ (115)
@@
$f_{U}(x)=\begin{cases}f(x),&d(x,U)<d(x,V),\\ \frac{1}{2}f(x),&d(x,U)=d(x,V),\\ 0,&d(x,U)>d(x,V),\end{cases}$ (116)
@@
To show exactness at ${{\mathfrak{g}}_{al}}(U)\oplus{{\mathfrak{g}}_{al}}(V)$, suppose $f\in{{\mathfrak{g}}_{al}}(U)\cap{{\mathfrak{g}}_{al}}(V)$. From the proof of Prop. 71, there exist $r\geq 0,C_{UV}>0$ such that $d(x,U^{r}\cap V^{r})\leq C_{UV}\max(d(x,U),d(x,V))$. We may assume that $C_{UV}\geq 1$, in which case for any $x\in{\mathbb{R}}^{n}$ and any $k\in{\mathbb{N}}$
$(1+d(x,U^{r}\cap V^{r}))^{k}\leq C^{k}_{UV}\left(1+\max(d(x,U),d(x,V))\right)^{k}.$ (117)
@@
$p_{k,U^{r}\cap V^{r}}(f)\leq C^{k}_{UV}\max(p_{k,U}(f),p_{k,V}(f)).$ (118)
Since by Corollary 72 one can represent $U\wedge V$ by $U^{r}\cap V^{r}$, we conclude that $f\in{{\mathfrak{g}}_{al}}(U\wedge V)$. ∎